Two trains are moving in the same direction at 80 km/hr and 30 km/hr. The faster train crosses a man in the slower train in 24 seconds. Find the length of the faster train.
Answer by mananth(16949)

Trains are moving in same direction
so effective speed = 80-30 = 50 km/h
Time taken to cross the man in slower train = 24 seconds= 24/3600 hours
The length of the faster train is the distance traveled.
let it be D
d = speed * time
d= 50 *24/3600
d= 24/72
d= 1/3 km
= 333 meters ( rounded off)